OpenStudy (anonymous):

Divide the following polynomial (a ^2^n - a^n - 6) (a^n + 8)

OpenStudy (kc_kennylau):

(a^2)^n or a^2^n?

OpenStudy (kc_kennylau):

Let x=a^n

OpenStudy (anonymous):

its a^2n (the 2n is together)

OpenStudy (kc_kennylau):

Let x=a^n

OpenStudy (anonymous):

How do i do that

OpenStudy (kc_kennylau):

a^2n = x^2

OpenStudy (anonymous):

how would i write that as the answer

OpenStudy (kc_kennylau):

substitute

OpenStudy (anonymous):

my answer is a^n - a^n -3/a^n+8